# SafePTR[Uncaptioned image]: Token-Level Jailbreak Defense in Multimodal LLMs via Prune-then-Restore Mechanism

### v1

@@ 4.4 Efficiency Evaluation Results
Training cost and Inference Time.
An ideal defense should be both effective and efficient. To assess this, we compare SafePTR with prior methods in training size (samples in K) and inference latency (secs), as shown in Tab. 7. SafePTR achieves leading performance on safety and utility benchmarks while remaining lightweight. It requires no training (vs. 1223K/ 71k for TGA/Immune) and runs in 3.51 secs with a one-pass pipeline, faster than multi-pass methods like ECSO or CoCA. These results underscore SafePTR’s practicality for real-world deployment, offering strong robustness and generalization without additional training or runtime cost.
@@
Table 6: Utility scores on MM-Vet. We report the accuracy across 7 capability categories. The best results are highlighted in bold.
@@
Table 9: Ablation study of proposed components. Performance of Harmful Token Pruning (HTP) and Benign Feature Restoration (BFR) on FigStep (FIG), MM-SafetyBench (MSB), MM-Vet, and MME.
@@ 5 Related Work
Jailbreak Attacks on Multimodal Large Language Models. Recent studies have demonstrated that MLLMs are highly susceptible to jailbreak attacks. Gong et al. (2023) transforms harmful text into typographic images to bypass safety filters, while Liu et al. (2023c) exploit query-relevant images for similar effect. Dong et al. (2023) craft visual adversarial examples to circumvent guardrails. Luo et al. (2024b) introduce JailBreakV-28K to benchmark attack transferability. However, these studies rarely analyze how jailbreak bypass safeguards, leaving internal vulnerability mechanisms of MLLMs largely unexplored. To bridge this gap, we investigate where, how, and which harmful multimodal tokens trigger jailbreaks, which directly informs our defense design.
Defenses on Multimodal Large Language Models. To counter adversarial threats, recent defenses for MLLMs have explored input transformation Yan et al. (2024), inference-time risk evaluation Zhang et al. (2024a), and prompt-based mitigation strategies Wang et al. (2024b). Other methods focus on hidden state monitoring and cross-modal safety transfer Liu et al. (2024c), inference-time alignment Wang et al. (2024a), and LLM-guided constitutional calibration Li et al. (2024). Despite these advancements, few works examine jailbreaks at the representation level, leaving models vulnerable to text-driven attacks and prone to overdefensive behavior. Guided by our finding that semantic drift from safety-aligned representations is a key failure factor, we propose SafePTR, an interpretable and training-free defense that preserves utility while enhancing robustness.

### v2

@@ 4.4 Efficiency Evaluation Results
Training cost and Inference Time. An ideal defense should be both effective and efficient. To assess this, we compare SafePTR with prior methods in training size (samples in K) and inference latency (secs), as shown in Tab. 7. SafePTR achieves leading performance on safety and utility benchmarks while remaining lightweight. It requires no training (vs. 1223K/ 71k for TGA/Immune) and runs in 3.51 secs with a one-pass pipeline, faster than multi-pass methods like ECSO or CoCA. These results underscore SafePTR’s practicality for real-world deployment, offering strong robustness and generalization without additional training or runtime cost.
@@
Table 6: Utility scores on $M$-Vet. We report the accuracy across 7 capability categories. The best results are highlighted in bold.
@@
Table 9: Ablation study of proposed components. Performance of Harmful Token Pruning (HTP) and Benign Feature Restoration (BFR) on FigStep (FIG), $M$-SafetyBench (MSB), $M$-Vet, and MME.
@@ 5 Related Work
Jailbreak Attacks on Multimodal Large Language Models. Recent studies have demonstrated that MLLMs are highly susceptible to jailbreak attacks. (gong2023figstep) transforms harmful text into typographic images to bypass safety filters, while (liu2023query) exploit query-relevant images for similar effect. (dong2023visual) craft visual adversarial examples to circumvent guardrails. (luo2024jailbreakv28k) introduce JailBreakV-28K to benchmark attack transferability. However, these studies rarely analyze how jailbreak bypass safeguards, leaving internal vulnerability mechanisms of MLLMs largely unexplored. To bridge this gap, we investigate where, how, and which harmful multimodal tokens trigger jailbreaks, which directly informs our defense design.
Defenses on Multimodal Large Language Models. To counter adversarial threats, recent defenses for MLLMs have explored input transformation (yan2024eyes), inference-time risk evaluation (zhang2024eta), and prompt-based mitigation strategies (wang2024adashield). Other methods focus on hidden state monitoring and cross-modal safety transfer (liu2024crossmodal), inference-time alignment (wang2024immune), and LLM-guided constitutional calibration (li2024coca). Despite these advancements, few works examine jailbreaks at the representation level, leaving models vulnerable to text-driven attacks and prone to overdefensive behavior. Guided by our finding that semantic drift from safety-aligned representations is a key failure factor, we propose SafePTR, an interpretable and training-free defense that preserves utility while enhancing robustness.
@@ 6 Conclusion
7 Acknowledgements
This study is supported by grants from the National Natural Science Foundation of China (Grant No. U23A20315, No. 62425208, No. U22A2097, No. 62122018, No. 62020106008), Shenzhen Science and Technology Program (No.JCYJ20240813114208012), Fundamental Research Funds for the Central Universities, and Natural Science Foundation of Sichuan Province (Grant No. 2025ZNSFSC1463).